OpenStudy (anonymous):

the slope of a line is -1/3, and the y-intercept is 10/3. What is the equation of the line written in general form?

OpenStudy (anonymous):

10x + 3y - 1 = 0?

OpenStudy (agent47):

y-10/3=(-1/3)(x-0) is the equation.. now general form..

OpenStudy (anonymous):

x+3y-10=0?

OpenStudy (agent47):

3y-10=-x x+3y=10

OpenStudy (anonymous):

Oh! ok! lol now I get it.

OpenStudy (agent47):

yea, you're right, but different sources are telling me diff things.. Depends on what your teacher told u the general form is.

OpenStudy (anonymous):

Oh ok...that makes sense. Thank you!